# OpenMonoGS-SLAM: Monocular Gaussian Splatting SLAM with Open-set Semantics

### v1

@@ IV EXPERIMENTS
MonoGS [matsuki2024gaussian] RGB PSNR ↑ 27.77 26.23 25.28 27.87 31.28 33.61 23.78 27.97 26.15
@@
Photo-SLAM [hhuang2024photoslam] RGB PSNR ↑ 30.37 29.73 26.65 31.97 35.27 28.74 30.72 29.25 30.62
@@
SEGS-SLAM [wen2025scaffold] RGB PSNR ↑ 33.54 31.25 27.27 35.09 38.56 38.64 34.42 33.90 29.35
@@
Hier-SLAM [li2024hier] RGB-D PSNR ↑ 33.57 30.16 31.15 33.46 37.36 39.12 31.27 30.83 33.19
@@ IV-A Experiments Setup
Implementation Details. Our OpenMonoGS-SLAM system is built upon the MASt3R-SLAM framework. We initialize the 3D Gaussian attributes from dense 3D point maps generated from the pretrained MASt3R and adopt the gsplat framework [ye2025gsplat] for efficient and scalable differentiable rendering of the 3D scene.
For multi-scale learning, we use $S=4$ scale levels and the memory threshold is set to $\tau_{m}=0.9$. The loss weight parameters are $\lambda_{\text{photo}}=1.0$, $\lambda_{\text{corr}}=0.05$, and $\lambda_{\text{lang}}=0.05$. For visual foundation models (VFMs), we utilize MASt3R, SAM (ViT-H), and CLIP (ViT-B) with their default configurations. All experiments are conducted on a single NVIDIA RTX 4090 GPU with 24GB VRAM. Mapping is performed for 30K iterations. Following MASt3R-SLAM’s preprocessing pipeline, we resize the image width to 512 pixels for both our model and baselines for evaluation.
To address open-set semantic segmentation, we incorporate Grounded-SAM [ren2024grounded], which enables segmentation via text prompt guidance without requiring predefined class labels. For fair comparisons with baseline methods, all offline models are trained using a consistent setting—selecting every 10th frame as a training view and optimizing the models for 30K iterations.
Dataset and Evaluation Metrics. We evaluate our method on the Replica [straub2019replica] dataset, which contains ground-truth semantic labels provided by [zhi2021place]. We additionally evaluate on the TUM RGB-D benchmark [tum] under a monocular setting, reporting tracking and reconstruction quality. For camera tracking evaluation, we compute the absolute trajectory error (ATE) using the root mean square error (RMSE) metric, implemented via the EVO [grupp2017evo] toolbox. To assess mapping quality, we evaluate photometric reconstruction using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure [wang2004ssim] (SSIM), and Learned Perceptual Image Patch Similarity [zhang2018lpips] (LPIPS). Semantic segmentation performance is quantified using mean Intersection over Union (mIoU). Frequency Weight IoU (FWIoU) and Accuracy (Acc) are employed for ablation studies.
Baselines. We evaluate tracking and novel view synthesis performance by comparing our method with representative monocular visual SLAM [matsuki2024gaussian, hhuang2024photoslam, wen2025scaffold] and RGB-D semantic SLAM [li2024dns, li2024sgs, li2024hier] baselines. For the evaluation of open-set segmentation, we evaluate our performance with offline feature-based segmentation methods [zhou2024feature, ye2024gaussian]. All baselines are implemented using their official code repositories. To ensure fairness, we fix the number of mapping iterations to 30k for all methods. If a baseline is originally trained with fewer iterations, we fine-tune it until it reaches 30k iterations. Furthermore, to maintain consistent input resolution, we resize all input images to a width of 512 pixels for both our method and the baselines.
@@ IV-B Results
Novel View Synthesis. We evaluate novel view rendering using all frames except the keyframes used for training. As shown in Tab. I and Fig. 2, our method consistently outperforms both visual and semantic SLAM baselines in most scenes, particularly in SSIM, indicating superior preservation of structural information. This improvement stems from jointly learning semantics and appearance; incorporating semantic cues helps maintain object-level consistency and preserve structural boundaries, leading to more accurate and coherent reconstructions than baselines. For fair evaluation in semantic SLAM comparisons, our model is trained using ground-truth semantic masks instead of SAM-generated masks. Note that while all semantic SLAM baselines utilize ground-truth depth maps as additional input, our method relies solely on RGB yet still achieves superior performance. In addition, Tab. III reports mapping performance on the TUM-D dataset, where our method achieves the highest PSNR and SSIM and the second-best LPIPS among all methods, demonstrating that the learned 3D Gaussian representation generalizes well beyond Replica and produces sharper, more perceptually faithful renderings than monocular baselines.
Camera Tracking. As shown in Tab. II, our method significantly outperforms existing monocular visual SLAM approaches in terms of absolute trajectory accuracy. Notably, it achieves strong robustness with low variance across a wide range of indoor scenes, consistently maintaining low ATE RMSE values. This stability can be attributed to the use of visual foundation models trained on large-scale, diverse datasets, which provide rich semantic and geometric priors. These priors enable our model to adopt effectively across the dataset and support more accurate and reliable camera pose estimation under challenging conditions. A similar trend is observed on the TUM-D sequences (Tab. III), where our method maintains competitive tracking accuracy on real-world indoor scenes under a monocular input setting.
@@
MonoGS 30.48 12.98 48.22 12.11 26.14 19.20 47.30 8.49 69.43
Photo-SLAM 10.63 1.60 22.56 3.36 2.11 30.02 6.95 2.04 16.38
SEGS-SLAM 9.25 1.07 35.09 0.22 1.70 0.74 0.62 1.11 33.40
@@
TABLE II: Camera tracking results on the Replica dataset. ATE RMSE in cm is reported. Our method demonstrates robust and consistent tracking performance across all scenes, achieving the lowest average error.
[…]
Visual SLAM MonoGS [matsuki2024gaussian] 4.40 20.88 0.694 0.358
Photo-SLAM [hhuang2024photoslam] 1.80 19.28 0.673 0.334
SEGS-SLAM [wen2025scaffold] 1.14 23.20 0.769 0.271
Semantic SLAM Ours (SAM1.0) 1.44 23.33 0.813 0.284
TABLE III: Quantitative comparisons on the TUM-D dataset.
Open-set Segmentation Results. To evaluate open-set segmentation performance using text prompts, we employ Grounded-SAM to generate segmentation masks conditioned on text prompts. Since both OpenMonoGS-SLAM and Feature 3DGS support prompt-based segmentation, we leverage the same evaluation text descriptions as prompts to obtain binary masks for each target concept. These binary masks are then applied to the rendered feature maps, enabling prompt-based segmentation within the visible regions of the scene. This setup allows us to fairly assess each model’s ability to localize and segment open-set concepts using only language-driven guidance, without relying on predefined semantic classes.
In contrast, Gaussian Grouping relies on a separate classifier for semantic prediction and does not support prompt-based segmentation directly on the rendered feature maps. As a result, we follow their official evaluation protocol to ensure a consistent comparison. Specifically, we compute the Intersection over Union (IoU) between the text-prompted binary masks generated by Grounded-SAM and the predicted segmentation masks produced by Gaussian Grouping. To account for open-set settings and reduce noise from irrelevant regions, we report the mean IoU only for cases where the overlap exceeds a threshold of 0.2, which aligns with the default threshold specified in their implementation.
As shown in Tab. IV, our method demonstrates superior open-set segmentation performance across all scenes. This improvement is largely attributed to our effective integration of CLIP-derived language features into the Gaussian semantic feature space using a memory-based mechanism. This facilitates more accurate alignment between visual features and a wide range of textual prompts, allowing the model to effectively recognize and segment diverse concepts (Fig. 3).
@@
Feature 3DGS 0.571 0.506 0.569 0.527 0.633 0.598 0.613 0.529 0.593
Gaussian Grouping 0.690 0.613 0.631 0.732 0.744 0.683 0.763 0.664 0.686
@@
TABLE IV: Quantitative comparisons of open-set segmentation on the Replica dataset. IoU from text prompts is reported.
Closed-set Segmentation Results. While our primary focus is open-set semantic SLAM, we additionally evaluate closed-set segmentation to demonstrate the effectiveness and general applicability of our framework using ground-truth semantic annotations.
Since all baselines directly utilize ground-truth semantic IDs during SLAM optimization, their predicted labels can be directly compared to ground-truth annotations using mean Intersection over Union (mIoU). In contrast, our method is inherently designed with a class-agnostic framework that operates without explicit semantic supervision. To facilitate a meaningful comparison, we introduce an additional loss term during SLAM optimization that incorporates ground-truth semantic IDs, allowing our model to align with the closed-set evaluation protocol. Specifically, we augment our existing self-supervised objectives with a cross-entropy loss to guide semantic label learning. After SLAM optimization, we follow the same evaluation protocol as the baselines by computing mIoU between predicted and ground-truth masks for each semantic ID.
As shown in Tab. V, our method achieves the best performance across all scenes except R1 and Of1, where it remains highly competitive. This demonstrates that, despite being tailored for open-set scenarios, our approach generalizes well to closed-set segmentation tasks.
@@
DNS-SLAM 0.742 0.818 0.831 0.748 0.671 0.798 0.759 0.721 0.589
SGS-SLAM 0.866 0.862 0.887 0.865 0.886 0.927 0.854 0.782 0.866
Hier-SLAM 0.804 0.869 0.767 0.844 0.817 0.807 0.830 0.740 0.793
Ours(Prior) 0.896 0.886 0.881 0.900 0.904 0.915 0.909 0.875 0.901
TABLE V: Quantitative comparisons of close-set segmentation on the Replica dataset. mIoU based on ground-truth semantic labels is reported.
Ablations and Analysis. To validate the effectiveness of each component in our framework, we perform comprehensive ablation studies, including both qualitative and quantitative evaluations, as presented in Fig. 4 and Tab. VI. All methods are trained for 30K iterations using the full set of scenes from the Replica dataset for a fair comparison.
We begin by analyzing the contribution of each loss term through individual ablations. Removing the multi-view contrastive loss (“w/o contrastive loss”) results in a substantial drop in mIoU by approximately 26%, along with notable declines in FWIoU and pixel accuracy. This highlights the importance of enforcing semantic consistency across views through multi-view contrastive learning. Similarly, excluding the regression loss (“w/o regression loss”) also results in degraded performance, highlighting its essential role in guiding semantic alignment. The combination of both losses yields a synergistic effect that is critical for achieving robust and reliable open-set segmentation. As shown in the visual results in Fig. 4, removing either loss leads to inaccurate and semantically inconsistent masks. In particular, the absence of the contrastive loss produces spatially fragmented and less coherent segmentation outputs.
We also ablate the memory attention mechanism by removing the memory bank used to accumulate masked CLIP features. Without the memory bank (“w/o memory”), mIoU drops by 19%, and FWIoU and accuracy decline by around 10%, confirming the pivotal role of memory-based attention in semantic learning. This indicates that leveraging past CLIP embeddings across frames helps enforce multi-view consistency, in contrast to relying solely on the current frame’s CLIP feature. These results collectively verify the design choices of our framework and demonstrate that each component contributes significantly to the final performance.
@@
Ours (SAM1.0) 0.645 0.685 0.730 34.47
TABLE VI: Ablations. Our method achieves better semantic segmentation and reconstruction performance, demonstrating the importance of each component in our framework.
@@ V CONCLUSION
We introduced OpenMonoGS-SLAM, a novel semantic SLAM framework capable of open-set segmentation under a monocular setup. By integrating visual foundation models (VFMs) such as MASt3R, SAM, and CLIP, our system achieves robust generalization across diverse tasks, enabling accurate monocular camera tracking and mapping, while providing rich semantic understanding in open-world environments. Furthermore, our method operates without any depth input or 3D semantic ground truth, relying entirely on self-supervised learning objectives to guide both geometric reconstruction and semantic understanding. Experimental results demonstrate that OpenMonoGS-SLAM achieves superior performance compared to existing baselines, highlighting the promise of integrating VFMs into SLAM. However, a current limitation lies in the lack of robustness to dynamic scenes, inherited from the MASt3R backbone. This opens future directions for leveraging models like [zhang2024monst3r], which are trained on dynamic environments, to enable even more general and wild-scene applicability.

### v2

@@ IV EXPERIMENTS
MonoGS [19] RGB PSNR ↑ 27.77 26.23 25.28 27.87 31.28 33.61 23.78 27.97 26.15
@@
Photo-SLAM [9] RGB PSNR ↑ 30.37 29.73 26.65 31.97 35.27 28.74 30.72 29.25 30.62
@@
SEGS-SLAM [33] RGB PSNR ↑ 33.54 31.25 27.27 35.09 38.56 38.64 34.42 33.90 29.35
@@
Hier-SLAM [16] RGB-D PSNR ↑ 33.57 30.16 31.15 33.46 37.36 39.12 31.27 30.83 33.19
@@ IV-A Experiments Setup
Implementation Details. Our OpenMonoGS-SLAM system is built upon the MASt3R-SLAM framework. We initialize the 3D Gaussian attributes from dense 3D point maps generated from the pretrained MASt3R and adopt the gsplat framework [37] for efficient and scalable differentiable rendering of the 3D scene.
For multi-scale learning, we use $S=4$ scale levels and the memory threshold is set to $\tau_{m}=0.9$. The loss weight parameters are $\lambda_{\text{photo}}=1.0$, $\lambda_{\text{corr}}=0.05$, and $\lambda_{\text{lang}}=0.05$. For visual foundation models (VFMs), we utilize MASt3R, SAM (ViT-H), and CLIP (ViT-B) with their default configurations. All experiments are conducted on a single NVIDIA RTX 4090 GPU with 24GB VRAM. Mapping is performed for 30K iterations, adding a mapping frame every 10 frames in periods where no keyframe is selected. Following MASt3R-SLAM’s preprocessing pipeline, we resize all input images to a width of 512 pixels for both our method and the baselines to ensure a fair comparison.
To address open-set semantic segmentation, we incorporate Grounded-SAM [24], which enables segmentation via text prompt guidance without requiring predefined class labels. For fair comparison with baselines, all offline models are trained using a consistent setting, selecting every 10th frame as a training view and optimizing for 30K iterations.
Dataset and Evaluation Metrics. We evaluate our method on the Replica [28] dataset, which contains ground-truth semantic labels provided by [40]. We also evaluate on the TUM RGB-D [29] and ScanNet [3] benchmark under a monocular setting, reporting tracking and reconstruction quality. For camera tracking evaluation, we compute the absolute trajectory error (ATE) using the root mean square error (RMSE) metric To assess mapping quality, we evaluate photometric reconstruction using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure [31] (SSIM), and Learned Perceptual Image Patch Similarity [39] (LPIPS). Semantic segmentation performance is quantified using mean Intersection over Union (mIoU). Frequency Weight IoU (FWIoU) and Accuracy (Acc) are used for ablation studies.
Baselines. We evaluate tracking and novel view synthesis performance by comparing our method with representative monocular visual SLAM [19, 9, 33, 25] and RGB-D semantic SLAM [17, 18, 16] baselines. For the evaluation of open-set segmentation, we evaluate our performance with offline feature-based segmentation methods [41, 36]. All baselines are implemented using their official code repositories. To ensure fairness, we fix the number of mapping iterations to 30k for all methods. If a baseline is originally trained with fewer iterations, we fine-tune it until it reaches 30k iterations.
@@ IV-B Results
Novel View Synthesis. We evaluate novel view rendering using all frames except the keyframes used for training. As shown in Tab. I and Fig. 2, our method consistently outperforms both visual and semantic SLAM baselines in most scenes, particularly in SSIM, indicating superior preservation of structural information. This improvement stems from jointly learning semantics and appearance; incorporating semantic cues helps maintain object-level consistency and preserve structural boundaries, leading to more accurate and coherent reconstructions. For fair evaluation in semantic SLAM comparisons, our model is trained using ground-truth semantic masks instead of SAM-based masks. Note that while semantic SLAM baselines utilize ground-truth depth as additional input, our method relies solely on RGB yet still achieves superior performance. We additionally report results on two real-world benchmarks. On TUM-D (Tab. II), our method achieves the highest PSNR and SSIM and the second-best LPIPS, demonstrating strong generalization beyond Replica. On ScanNet (Tab. III), our method attains the best rendering quality across all metrics against monocular baselines. Moreover, despite using SAM-based masks rather than ground-truth annotations, it outperforms all semantic SLAM baselines across all metrics.
[…]
Visual SLAM MonoGS [19] 4.40 20.88 0.694 0.358
Photo-SLAM [9] 1.80 19.28 0.673 0.334
SEGS-SLAM [33] 1.14 23.20 0.769 0.271
Semantic SLAM Ours (SAM1.0) 1.44 23.33 0.813 0.284
TABLE II: Quantitative comparisons on the TUM-D dataset.
[…]
Visual SLAM Splat-SLAM [25] 8.25 24.12 0.796 0.359
SEGS-SLAM [33] 7.66 23.62 0.846 0.304
Semantic SLAM SGS-SLAM [18] 9.86 22.27 0.753 0.348
Hier-SLAM [16] 10.53 21.78 0.739 0.355
Ours (SAM1.0) 5.39 26.21 0.860 0.273
TABLE III: Quantitative comparisons on the ScanNet dataset.
Camera Tracking. As shown in Tab. IV, our method significantly outperforms existing monocular visual SLAM approaches in absolute trajectory accuracy. It is also robust across diverse indoor scenes, exhibiting low variance and consistently achieving low ATE RMSE. We attribute this stability to visual foundation models trained on large-scale and diverse data, which provide rich semantic and geometric priors and thus enable more reliable pose estimation under challenging conditions. On the real-world TUM-D benchmark (Tab. II), our method attains the second-lowest ATE with a small gap to the best baseline, indicating competitive tracking performance under monocular input, while on ScanNet (Tab. III) it achieves the lowest ATE among all baselines, demonstrating strong robustness on real-world indoor scenes.
@@
MonoGS [19] 30.48 12.98 48.22 12.11 26.14 19.20 47.30 8.49 69.43
Photo-SLAM [9] 10.63 1.60 22.56 3.36 2.11 30.02 6.95 2.04 16.38
SEGS-SLAM [33] 9.25 1.07 35.09 0.22 1.70 0.74 0.62 1.11 33.40
@@
TABLE IV: Camera tracking results on the Replica dataset. ATE RMSE in cm is reported. Our method demonstrates robust and consistent tracking performance across all scenes, achieving the lowest average error.
Open-set Segmentation Results. To evaluate open-set segmentation with text prompts, we use Grounded-SAM [24] to generate text-conditioned 2D masks. Since both OpenMonoGS-SLAM and Feature 3DGS [41] support prompt-based segmentation on rendered feature maps, we use the same evaluation text descriptions as prompts to obtain binary masks for each target concept. These masks are applied to the rendered feature maps to produce prompt-based segmentations within the visible regions, enabling a fair assessment of language-driven concept localization without relying on predefined semantic classes.
In contrast, Gaussian Grouping [36] predicts semantics via a separate classifier and does not support prompt-based segmentation on rendered features. Therefore, following the official Gaussian Grouping protocol, we compute IoU between the text-prompted masks generated by Grounded-SAM and the predicted masks from Gaussian Grouping. To reduce the effect of irrelevant regions in the open-set setting, we report mean IoU only for pairs with IoU >0.2>0.2, matching the default threshold in their implementation.
As shown in Tab. V, our method achieves the best open-set segmentation performance across all scenes. This improvement is largely attributed to our effective memory-based integration of CLIP language features into the Gaussian feature space, which improves alignment between rendered features and diverse textual prompts, leading to more accurate open-set segmentation (Fig. 3).
@@
Feature 3DGS[41] 0.571 0.506 0.569 0.527 0.633 0.598 0.613 0.529 0.593
Gaussian Grouping[36] 0.690 0.613 0.631 0.732 0.744 0.683 0.763 0.664 0.686
@@
TABLE V: Quantitative comparisons of open-set segmentation on the Replica dataset. mIoU from text prompts is reported.
Closed-set Segmentation Results. While our primary focus is open-set semantic SLAM, we additionally evaluate closed-set segmentation using ground-truth semantic annotations to demonstrate the general applicability of our framework.
All semantic SLAM baselines directly use ground-truth semantic IDs during optimization, enabling straightforward evaluation by computing mean Intersection over Union (mIoU) between predicted and ground-truth labels. In contrast, our method is class-agnostic and does not assume explicit semantic supervision. To enable a fair closed-set comparison, we add an auxiliary cross-entropy loss term during SLAM optimization that leverages the ground-truth semantic IDs, while keeping the rest of our objectives unchanged. After optimization, we follow the same protocol as the baselines and report mIoU between the predicted and ground-truth masks for each semantic class.
As shown in Tab. VI, our method achieves the best performance across all scenes except R1 and Of1, where it remains highly competitive. These results indicate that, although designed for open-set scenarios, the proposed approach also effectively transfers to closed-set segmentation.
@@
DNS-SLAM[17] 0.742 0.818 0.831 0.748 0.671 0.798 0.759 0.721 0.589
SGS-SLAM[18] 0.866 0.862 0.887 0.865 0.886 0.927 0.854 0.782 0.866
Hier-SLAM[16] 0.804 0.869 0.767 0.844 0.817 0.807 0.830 0.740 0.793
Ours(GT) 0.896 0.886 0.881 0.900 0.904 0.915 0.909 0.875 0.901
TABLE VI: Quantitative comparisons of close-set segmentation on the Replica dataset. mIoU based on ground-truth semantic labels is reported.
Ablations and Analysis. To validate the effectiveness of each component in our framework, we conduct comprehensive ablation studies with both qualitative and quantitative evaluations as presented in Fig. 4, Fig. 5 and Tab. VII. All methods are trained for 30K iterations on the full set of Replica scenes for a fair comparison. For segmentation evaluation, we use all object masks available in the semantic ground truth as prompts for each scene.
We first analyze the contribution of each loss term. Removing the multi-view contrastive loss (“w/o contrastive loss”) causes a substantial mIoU drop of approximately 27%, accompanied by clear declines in FWIoU and pixel accuracy, highlighting the importance of enforcing cross-view semantic consistency. Excluding the regression loss (“w/o regression loss”) also degrades performance, indicating its essential role in guiding semantic alignment. As shown in Fig. 5, removing either loss leads to inaccurate and semantically inconsistent masks. In particular, without the contrastive loss, predictions become spatially fragmented and less coherent.
Figure 4: Qualitative comparison of multi-scale ablations on the Replica dataset. Multi-scale supervision produces more coherent masks with finer details.
We further ablate the proposed multi-scale strategy by training with only a single scale level (“only coarse scale” or “only fine scale”). As reported in Tab. VII, using a single scale consistently underperforms the full model, indicating that multi-scale supervision is crucial for handling objects with diverse sizes and granularities. Qualitative results in Fig. 4 show that the coarse-only setting tends to over-group regions and inaccurate boundaries, whereas the fine-only setting often produces noisier and less stable masks. In contrast, combining multiple scales yields more coherent object-level segmentation while preserving fine details.
We also ablate the memory attention mechanism by removing the memory bank that accumulates masked CLIP features. Without the memory bank (“w/o memory”), mIoU drops by 21%, and both FWIoU and accuracy decrease by 10%, confirming the pivotal role of memory-based attention in semantic learning. This suggests that aggregating CLIP embeddings across frames provides more consistent semantic guidance than relying only on the current frame. Overall, these results verify our design choices and show that each component contributes significantly to the final performance.
@@
only coarse scale 0.623 0.704 0.746 33.84
only fine scale 0.647 0.696 0.743 33.54
@@
Ours (SAM1.0) 0.660 0.708 0.754 34.47
TABLE VII: Ablations. Our method achieves better semantic segmentation and reconstruction performance, demonstrating the importance of each component in our framework.
@@ V CONCLUSION
We introduced OpenMonoGS-SLAM, a novel semantic SLAM framework capable of open-set segmentation under a monocular setup. By integrating visual foundation models (VFMs) such as MASt3R, SAM, and CLIP, our system achieves robust generalization across diverse tasks, enabling accurate monocular camera tracking and mapping, while providing rich semantic understanding in open-world environments. Furthermore, our method operates without any depth input or 3D semantic ground truth, relying entirely on self-supervised learning objectives to guide both geometric reconstruction and semantic understanding. Experimental results demonstrate that OpenMonoGS-SLAM achieves superior performance compared to existing baselines, highlighting the promise of integrating VFMs into SLAM. However, a current limitation lies in the lack of robustness to dynamic scenes, inherited from the MASt3R backbone. This opens future directions for leveraging models like [38], which are trained on dynamic environments, to enable even more general and wild-scene applicability.